OpenStudy (yellowlegoguy99):

A square has side lengths that measure 8 inches. A rectangle has a length of ten inches and a width of 5 inches. How many more square inches are in the area of the square than the area of the rectangle? 1 square inch 2 square inches 7 square inches fourteen square inches

OpenStudy (yellowlegoguy99):

@e.cociuba

OpenStudy (yellowlegoguy99):

@halsdkjfhakjs

OpenStudy (anonymous):

First figure out the area of the square and the rectangle. You should be able to do that.

OpenStudy (yellowlegoguy99):

Area of square = 64 Inches

OpenStudy (anonymous):

Yes, perfect.

OpenStudy (yellowlegoguy99):

Area of rectangle is 50 inches

OpenStudy (anonymous):

@e.cociuba dont even bother replying, i got this under control

OpenStudy (e.cociuba):

Area of Square: 8^2=64 square inches Area of Rectangle: 10x5=50 square inches Difference: 64-50=14 sq in

OpenStudy (anonymous):

yes

OpenStudy (yellowlegoguy99):

64 - 50 = 14 square inches

OpenStudy (anonymous):

Sure, just give him the answer, I am actually trying to teach him something.

OpenStudy (e.cociuba):

@halsdkjfhakjs finally something you can do

OpenStudy (yellowlegoguy99):

Lol

OpenStudy (anonymous):

:)

OpenStudy (yellowlegoguy99):

Who should I give the medal to...?

OpenStudy (anonymous):

And yes yellow, that is correct, I'm glad you figured it out by yourself.

OpenStudy (e.cociuba):

Eh, give it to @halsdkjfhakjs I don't help people for medals. Useless.

OpenStudy (yellowlegoguy99):

Wait a sec..I have more questions

OpenStudy (anonymous):

Okay, happy to help.

OpenStudy (yellowlegoguy99):

What is the perimeter of this parallelogram? A parallelogram is shown with the bases each 8 inches, the height 2 inches, and the length of one of the sides 3 inches thirteen inches sixteen inches twenty-two inches twenty-four inches

OpenStudy (yellowlegoguy99):

Here is the image

OpenStudy (anonymous):

Well perimeter is just all the sides added together.

OpenStudy (anonymous):

So, just add them all together and tell me what you get.

OpenStudy (yellowlegoguy99):

24

OpenStudy (anonymous):

Yes, that is the right answer.

OpenStudy (yellowlegoguy99):

Is that right?

OpenStudy (anonymous):

Yes, I just said.

OpenStudy (yellowlegoguy99):

What is the perimeter of this isosceles triangle? An isosceles triangle with base thirty-six meters, two sides both equal to thirty meters, and height twenty-four meters ninety meters ninety-six meters one hundred ninety-two meters four hundred thirty-two meters

OpenStudy (yellowlegoguy99):

Here is the pic

OpenStudy (anonymous):

Sorry, I have to go now.

OpenStudy (anonymous):

@e.cociuba Could you help yeloowguy99 from here on out? Thanks

OpenStudy (yellowlegoguy99):

@e.cociuba

OpenStudy (anonymous):

@halsdkjfhakjs and @e.cociuba stop arguing u both helped him in different ways